How do you solve this system of linear equations 4a=7-b and 4a-3-b using the substitution method
Answer by stanbon(75887)

solve this system of linear equations
4a=7-b and 4a= 3-b using the substitution method.
---------
Since both equal 4a,
7-b = 3-b
----
That is a contradiction.
Conclusion: No solution
==========================
Cheers,
Stan H.